in the problem h(x)=x^2+2x-3, i know the domain is all real numbers, but how come the range is [-4,to infinity). how do I find the range of the problem?
Answer by stanbon(75887)

in the problem h(x)=x^2+2x-3, i know the domain is all real numbers, but how come the range is [-4,to infinity). how do I find the range of the problem?
----
Since the vertex is either at the bottom or at the top of the range,
find the vertex.
----
x = -b/(2a) = -2/(2*1) = -1
---
f(-1) = (-1)^2 +2(-1)-3 = 1-2-3 = -4
----
Since "a" is positive the parabola opens up,
so -4 is the lowest y-value.
Range: [-4,+oo)
======================
Cheers,
Stan H.